OpenStudy (horsegal244):

What is the simplified expression for 3^3•3^3/3^4?

OpenStudy (horsegal244):

@Michele_Laino

OpenStudy (horsegal244):

a. 3^0 b. 3^1 c. 3^2 d. 3^4

OpenStudy (thread252):

the answer is 9

OpenStudy (horsegal244):

so C

OpenStudy (thread252):

yup

OpenStudy (seascorpion1):

I'm here

OpenStudy (michele_laino):

If I apply the rules of multiplication and division between powers with the same base, I get: \[\huge \frac{{{3^3} \cdot {3^3}}}{{{3^4}}} = {3^{3 + 3 - 4}} = ...?\]